# Effect of Domain Generalization Techniques in Low-Resource Systems

### v1

@@ Abstract
Machine learning models typically assume that training and test data follow the same distribution, an assumption that often fails in real-world scenarios due to distribution shifts. This issue is especially pronounced in low-resource settings, where data scarcity and limited domain diversity hinder robust generalization. Domain generalization (DG) approaches address this challenge by learning features that remain invariant across domains, often using causal mechanisms to improve model robustness. In this study, we examine two distinct causal DG techniques in low-resource natural language tasks. First, we investigate a causal data augmentation (CDA) approach that automatically generates counterfactual examples to improve robustness to spurious correlations. We apply this method to sentiment classification on the NaijaSenti Twitter corpus, expanding the training data with semantically equivalent paraphrases to simulate controlled distribution shifts. Second, we explore an invariant causal representation learning (ICRL) approach using the DINER framework, originally proposed for debiasing aspect-based sentiment analysis. We adapt DINER to a multilingual setting by constructing Afri-SemEval, a new benchmark derived from the SemEval-2014 Task 4 dataset via translation, covering 17 African languages. Through comprehensive experiments, we compare the effectiveness of these techniques in mitigating domain shifts and improving model performance. Our findings demonstrate that both approaches enhance robustness to unseen domains: counterfactual data augmentation yields consistent cross-domain accuracy gains in sentiment classification, while causal representation learning with DINER improves out-of-distribution performance in multilingual sentiment analysis, albeit with varying gains across languages. These results highlight the complementary strengths and trade-offs of data-level vs. representation-level causal interventions in low-resource AI settings.
@@ 1 Introduction
Machine learning (ML) models have achieved remarkable success in diverse natural language processing (NLP) applications, yet their effectiveness largely depends on the assumption that training and test data are independently and identically distributed (i.i.d.). In practice, data distributions often shift across time, geography, and linguistic context, leading to significant performance degradation on unseen domains. This problem is especially acute in low-resource settings, where data scarcity and linguistic variation limit the ability of models to generalize robustly[1]
Domain generalization (DG) seeks to mitigate this challenge by learning representations that remain invariant across multiple source domains without direct access to the target distribution. Within DG, causal inference provides a principled foundation: by focusing on stable cause–and–effect relationships, models can identify invariant mechanisms that hold under interventions or domain shifts[2]. Causal modeling aims to capture the underlying generative processes common to all domains, thereby improving robustness when faced with distributional changes.
In this study, we examine two complementary causal DG approaches for low-resource sentiment analysis in African languages:
  • Invariance via Causal Data Augmentation (CDA): We leverage GPT-4o-mini to generate semantically equivalent paraphrases that preserve sentiment polarity while introducing lexical and syntactic diversity. This process expands the training data and simulates a controlled domain shift (from original domain $D_{1}$ to augmented domain $D_{2}$) under the constraint $P(y\mid x^{\prime})=P(y\mid x)$. Using the Yoruba and Igbo subsets of the NaijaSenti dataset[3], we assess how exposure to paraphrastic variation improves generalization and reduces the invariance gap $\Delta_{\text{inv}}$.
  • Invariance via Causal Representation Learning (CRL): We adopt the DINER framework[4], which models sentiment prediction as a Structural Causal Model (SCM) with explicit backdoor adjustments and counterfactual reasoning to isolate causal sentiment paths. To extend DINER to low-resource African contexts, we introduce Afri-SemEval, a multilingual benchmark translated from the SemEval-2014 Task 4 ABSA dataset[5], covering 17 African languages. Each translated instance retains the original review, target aspect, and sentiment polarity, allowing us to evaluate model invariance across a diverse set of languages and linguistic conditions.
By integrating these two frameworks, CDA for data-level causal perturbations and CRL for representation-level causal learning, we evaluate whether causal interventions at both levels can jointly enhance robustness under distributional shifts. Empirical results across Yoruba, Igbo, and the 17 Afri-SemEval languages show consistent gains in cross-domain accuracy and reduced sensitivity to spurious correlations. In particular, augmenting training data with counterfactual paraphrases yields more invariant performance between original and shifted distributions, while causal representation learning improves out-of-distribution generalization in multilingual sentiment analysis, with the magnitude of improvement varying based on linguistic and data characteristics.
Our contributions are threefold:
  • We propose a unified experimental design that combines data-level and representation-level causal interventions for low-resource African sentiment analysis.
  • We introduce Afri-SemEval, a multilingual extension of the SemEval-2014 ABSA dataset (translated into 17 African languages) to facilitate research on domain generalization in multilingual contexts.
  • We empirically demonstrate that incorporating causal mechanisms, through both data augmentation and causal representation learning, enhances generalization and robustness under distribution shifts in multilingual, low-resource NLP tasks.
2 Related Works
2.1 Domain Generalization and the Low-Resource Paradox
Recent comprehensive surveys reveal fundamental challenges in achieving OOD robustness for NLP systems [6, 7]. While domain adaptation techniques assume access to unlabeled target data, domain generalization confronts a more stringent constraint: models must extrapolate to completely unseen distributions using only source domain knowledge. Critically, the DomainBed evaluation demonstrated that carefully tuned empirical risk minimization often matches or exceeds specialized DG algorithms [8], suggesting that algorithmic complexity may not always translate to practical gains—a finding particularly relevant when computational resources are constrained.
For African languages, this theoretical challenge becomes acute. The NaijaSenti corpus [3] and AfriSenti benchmark [9] represent pioneering efforts, yet even these datasets contain only thousands of examples per language compared to millions available for high-resource languages. Community-driven documentation reveals that 88% of African languages lack basic computational resources [10], creating what we term a “low-resource paradox”: the very languages requiring the most robust generalization techniques have the least data to learn invariant representations from. Analysis of counterfactual data reveals that spurious correlations in small datasets can achieve training accuracies comparable to true causal features [11], explaining why models trained on limited data catastrophically fail when deployment conditions deviate even slightly from training distributions.
2.2 From Correlation to Causation: Theoretical Foundations
The causal revolution in machine learning offers a principled solution: if models learn causal mechanisms rather than statistical associations, they should maintain performance under distribution shifts [12, 13]. Comprehensive taxonomies of causal inference applications in NLP systematically categorize text as treatment, outcome, or confounder [14], yet most work focuses on high-resource English datasets, leaving open questions about scalability to low-resource, morphologically complex languages.
Two paradigms have emerged for implementing causal invariance. Invariant Risk Minimization penalizes representations where optimal classifiers vary across training environments [15], though subsequent analyses revealed optimization challenges and limited guarantees beyond linear settings [8]. More recently, explicit intervention methods construct task-specific Structural Causal Models to surgically remove spurious pathways [7, 2]. Our work bridges these paradigms by evaluating both data-level counterfactual generation and model-level causal debiasing in a unified low-resource experimental framework.
2.3 Counterfactual Augmentation and LLM-Based Data Generation
Counterfactual data augmentation, where annotators make minimal edits to flip labels while preserving coherence, demonstrates that models trained on counterfactually-augmented data exhibit 40-60% reduction in sensitivity to spurious features [11]. However, human annotation at scale remains prohibitive for low-resource languages. Active learning approaches with counterfactual generative models have achieved 18-20% robustness improvements using only 10% human-annotated counterfactuals [16]—a crucial finding for resource-constrained scenarios.
The emergence of instruction-tuned LLMs as data augmentation engines has been both promising and problematic. Recent experiments demonstrate that modern LLMs can expand 500 labeled examples to 5,000 synthetic samples with comparable downstream performance [17], but critically note that augmentation quality degrades substantially for low-resource languages like Tamil. This observation motivates our systematic evaluation of GPT-4o-mini [18] on Yoruba and Igbo—languages where paraphrase quality cannot be assumed but must be empirically validated. Unlike generic data augmentation techniques operating on token-level perturbations without semantic constraints [19], our approach explicitly enforces sentiment preservation through prompt engineering to maintain the counterfactual validity condition $P(y|x^{\prime})=P(y|x)$.
2.4 Causal Debiasing for Aspect-Based Sentiment Analysis
Aspect-Based Sentiment Analysis exposes fundamental weaknesses in neural architectures. The ARTS adversarial benchmark revealed that state-of-the-art ABSA models suffer up to 69.73% accuracy drops when aspect-sentiment correlations are manipulated [20], confirming that models exploit spurious statistical patterns [21, 22, 23] rather than genuine compositional understanding. While various debiasing strategies have been proposed [24, 25], most address single-variable inputs and fail to account for ABSA’s inherent multi-input structure.
The DINER framework represents a methodological advance by constructing an SCM distinguishing review-level confounding from aspect-level spurious correlation, applying backdoor adjustment to the former and counterfactual TIE estimation to the latter[4]. This hybrid strategy outperforms single-intervention baselines on English ABSA benchmarks. However, its applicability beyond English remains unexplored—a critical gap given that the SemEval-2014 Task 4 benchmark [5] exists only in English and a handful of European languages.
Our contribution addresses this gap through Afri-SemEval, translating the canonical ABSA benchmark into 17 typologically diverse African languages. This enables the first systematic evaluation of causal debiasing methods in multilingual low-resource contexts. Leveraging African-focused multilingual models with 17-language adaptive fine-tuning [26] and data-efficient pretraining [27], we investigate whether causal interventions transfer across linguistic typologies. The mixed results we observe illuminate a crucial limitation: translation-induced artifacts (word order changes, morphological mismatches, aspect-review misalignment) introduce unintended distribution shifts that interact with the intended cross-lingual domain shift [7]. This dual-shift phenomenon suggests that future causal DG methods must be robust not only to natural linguistic variation but also to the systematic noise inherent in cross-lingual transfer pipelines—a challenge distinct from the monolingual OOD scenarios dominating current literature.
@@ 3 Methodology
This section presents the methodological design adopted in this study, comprising two complementary experimental frameworks: (1)Invariance via Causal Data Augmentation and (2) Invariance via Causal Representation Learning.
@@ 3.1 Invariance via Causal Data Augmentation
This study investigates causal data augmentation as a strategy for improving sentiment classification in low-resource African languages through enhanced representation invariance. The approach is grounded in the principles of causal representation learning: if the sentiment label $y$ causally depends on the semantic content of a tweet $x$, then augmentations that perturb only non-causal linguistic features (e.g., surface form, style, or syntax) should preserve the conditional distribution $P(y|x)$. By systematically generating linguistically diverse yet semantically equivalent paraphrases, the model learns to focus on the underlying causal semantics rather than superficial correlations, leading to improved generalization across domains.
3.1.1 Data and Resources
@@
We utilize the Yoruba and Igbo subsets of the NaijaSenti corpus[3], a multilingual sentiment-labeled Twitter dataset covering four Nigerian languages: Yoruba, Igbo, Hausa, and Nigerian Pidgin. Each instance is a tweet–label pair $(x_{i},y_{i})$, where $x_{i}$ represents the tweet and $y_{i}\in\{\text{positive},\text{negative},\text{neutral}\}$ is its sentiment label.
3.1.2 Training Configuration
Let $\mathcal{D}_{1}=\{(x_{i},y_{i})\}_{i=1}^{N}$ denote the original data distribution drawn from $P(X,Y)$. We define an augmentation function $f:\mathcal{X}\rightarrow\mathcal{X}^{\prime}$ that generates paraphrased examples $x^{\prime}=f(x)$ such that the sentiment semantics are preserved, i.e.,
$P(y|x^{\prime})=P(y|x).$ (1)
To achieve this, we employ the GPT-4o-mini model[18] with a carefully designed prompt that instructs the model to generate paraphrases that preserve sentiment polarity while introducing lexical and syntactic diversity. The prompt template used is:
@@ 3.1.2 Training Configuration
"You have been provided with {language} tweet. Paraphrase ’{tweet}’ in a way that preserves its sentiment and meaning but changes its wording. Return only the paraphrased tweet."
This produces a new dataset $\mathcal{D}_{2}=\{(x^{\prime}_{i},y_{i})\}_{i=1}^{N}$ that shares semantic equivalence with $\mathcal{D}_{1}$ but differs in linguistic realization, thereby simulating a controlled domain shift:
$\begin{split}P_{\mathcal{D}_{2}}(X^{\prime})&\neq P_{\mathcal{D}_{1}}(X),\quad\text{while}\\ P_{\mathcal{D}_{2}}(Y|X^{\prime})&\approx P_{\mathcal{D}_{1}}(Y|X).\end{split}$ (2)
We train three distinct logistic regression classifiers for each language subset; (1)Original-only–trained on $\mathcal{D}_{1}$. (2) Paraphrased-only–trained on $\mathcal{D}_{2}$. (3) Combined–trained on $\mathcal{D}_{\text{comb}}=\mathcal{D}_{1}\cup\mathcal{D}_{2}$ (50-50 mix).
Each model is evaluated under a cross-domain testing scheme to measure generalization under distributional shift; Models trained on $\mathcal{D}_{1}$ are tested on $\mathcal{D}_{1}$, $\mathcal{D}_{2}$ and $\mathcal{D}_{\text{comb}}$, Models trained on $\mathcal{D}_{2}$ are tested on $\mathcal{D}_{2}$, $\mathcal{D}_{1}$ and $\mathcal{D}_{\text{comb}}$, Models trained on $\mathcal{D}_{\text{comb}}$ are tested on all three domains.
@@
For each configuration, the model $h_{\theta}$ minimizes the empirical cross-entropy loss:
$\hat{\mathcal{L}}(\theta)=\frac{1}{|\mathcal{D}_{\text{train}}|}\sum_{(x,y)\in\mathcal{D}_{\text{train}}}\ell(h_{\theta}(g_{\psi}(\phi(x))),y),$ (3)
where $\phi$ represents text preprocessing (tokenization, normalization, stopword removal), and $g_{\psi}$ denotes the embedding function implemented using the paraphrase-multilingual-MiniLM-L12-v2[28] sentence transformer.
0ur Model performance was evaluated using accuracy computed over multiple domains, and invariance was quantified by measuring performance consistency across test domains:
$\Delta_{\text{inv}}=|\text{Acc}_{\mathcal{D}_{1}}-\text{Acc}_{\mathcal{D}_{2}}|,$ (4)
where smaller $\Delta_{\text{inv}}$ indicates stronger invariance to paraphrastic shifts. This design was adopted to enable a principled test of the causal invariance hypothesis, which posits that models trained with causal data augmentation should minimize $\Delta_{\text{inv}}$ and achieve superior generalization in low-resource settings.

### v2

# When Distributions Shifts: Causal Generalization for Low-Resource Languages

@@ Abstract
Machine learning models often fail under distribution shifts, a problem exacerbated in low-resource settings where limited data restricts robust generalization. Domain generalization (DG) methods address this challenge by learning representations that remain invariant across domains, frequently leveraging causal principles. In this work, we study two causal DG approaches for low-resource natural language processing. First, we apply causal data augmentation using GPT-4o-mini to generate counterfactual paraphrases for sentiment classification on the NaijaSenti Twitter corpus in Yoruba and Igbo. Second, we investigate invariant causal representation learning with the Debiasing in Aspect Review (DINER) framework for aspect-based sentiment analysis. We extend DINER to a multilingual setting by introducing Afri-SemEval, a dataset of 17 languages translated from SemEval-2014 Task. Experiments show improved robustness to unseen domains, with consistent gains from counterfactual augmentation and enhanced out-of-distribution performance from causal representation learning across multiple languages.
@@ 1 Introduction
Machine learning models have achieved notable success in natural language processing, but their effectiveness relies on the assumption that training and test data are identically distributed. In practice, data distributions shift across time, geography, and linguistic contexts, leading to performance degradation on unseen domains. This challenge is severe in low-resource settings. In a study conducted by Adebara et al. [1], they introduced the “Sahara” benchmark for African NLP and showed that out of 517 African languages in their benchmark, only 45 languages were present in more than one dataset, with most appearing only in basic language identification tasks. Such extreme data sparsity and fragmentation make it infeasible to rely on scale alone, amplifying the impact of distribution shifts and the need for domain generalization methods that can learn invariant representations under limited and heterogeneous data conditions [2, 3, 4, 5].
In machine learning, domain generalization (DG) addresses the challenge of performing well on unseen domains by learning representations that remain invariant across multiple source domains without access to the target distribution. This DG formulation and its associated solutions also extend to low-resource settings, where data scarcity makes robustness to distribution shifts especially important. Recent DG approaches increasingly emphasize learning stable and invariant features across domains; causal perspectives have been shown to be one effective way to formalize and motivate such invariance under domain shifts [6].
In this study, we examine two complementary causal DG approaches for low-resource sentiment analysis: Causal Data Augmentation (CDA), which generates sentiment-preserving paraphrases to simulate controlled domain shifts, and Causal Representation Learning (CRL), which applies structural causal models with backdoor adjustment and counterfactual reasoning to isolate causal sentiment paths. Figure 1 illustrates our unified framework. For CDA, we use the Yoruba and Igbo subsets of NaijaSenti [7]. For CRL, we introduce Afri-SemEval, a new benchmark translating SemEval-2014 Task 4 [8] into 17 African languages.
Figure 1: Overview of our unified causal framework. Left (CDA): Paraphrased data is generated via GPT-4o-mini under the constraint $P(Y|X^{\prime})=P(Y|X)$, then combined with original data for invariant representation learning. Right (CRL): The DINER mechanism applies aspect-context alignment, backdoor adjustment, and counterfactual reasoning to produce debiased sentiment predictions.
Our contributions are: (1) a unified framework combining data-level and representation-level causal interventions for low-resource sentiment analysis; (2) Afri-SemEval, a multilingual ABSA benchmark covering 17 African languages; and (3) empirical evidence that causal mechanisms enhance generalization under distribution shifts in low-resource multilingual NLP.
In the rest of this paper, Section 2 reviews related work on domain generalization and causal NLP. Section 3 details our methodology for both CDA and CRL approaches. Section 4 presents experimental results and analysis, and Section 5 concludes the paper, discusses limitations and future directions.
2 Related Work
Domain generalization (DG) in NLP aims to improve model robustness under distribution shifts.
Table 1 summarizes key Domain Generalization approaches that have been explored by researchers, highlighting whether they use Causal Data Augmentation (CDA) or Causal Representation Learning (CRL) and their main contributions and limitations.
While prior work [9] have explored causal data augmentation (CDA) for low-resource languages, controlled sentiment-preserving paraphrasing for simulating causal domain shifts is largely unexplored. Additionally, causal representation learning (CRL) has mostly been applied to high-resource languages.
Our study explores two causal frameworks for low-resource African languages; CDA and CRL. Unlike previous works, we explored CDA via sentiment-preserving paraphrases in Yoruba and Igbo. For CRL, we introduced Afri-SemEval, a multilingual ABSA dataset covering 17 African languages, enabling evaluation of domain generalization under distribution shifts in low-resource multilingual settings.
Authors Technique Research Goal CDA/CRL Key Contribution
Wu et al. [10] Multilingual ABSA Benchmark (M-ABSA) Multilingual aspect-based sentiment analysis CRL Addressed data scarcity via benchmark construction and fine-tuning of generative models.
Muhammad et al. [7]; Muhammad et al. [11] Dataset Creation (NaijaSenti, AfriSenti) Sentiment analysis for African languages CDA Pioneered African language sentiment datasets but limited to thousands of examples.
Ibrahim et al. [12]; Maqsood et al. [13] Low-resource Benchmark Creation Sentiment analysis for Hausa and Urdu CRL Introduced language-specific benchmarks that improve coverage but not address robustness in domain shifts.
Kaushik et al. [14] Counterfactual Augmentation Reduce spurious feature sensitivity CDA Showed 40–60% reduction in spurious correlations via counterfactuals.
Šmíd et al. [15] Constrained Decoding Syntactic validity in cross-lingual generation CRL Enforced syntactic constraints during generation without addressing semantic or causal robustness.
Balashankar et al. [16] Active Learning + Counterfactual Generation Efficient robustness with minimal annotation CDA Achieved 18–20% robustness gains with 10% human annotation.
Møller et al. [9] GPT-4 Data Augmentation Low-resource classification tasks CDA Noted quality degradation for low-resource languages.
Wu et al. [17] DINER Debias ABSA via causal inference CRL Backdoor adjustment and counterfactual reasoning to English ABSA.
Table 1: Overview of domain generalization approaches relevant to low-resource NLP. Techniques are categorized by authors, method, research goal, and type of causal intervention: Causal Data Augmentation (CDA) or Causal Representation Learning (CRL). Key contributions highlight each method’s main achievements and limitations.
@@ 3 Methodology
This section presents the methodological design adopted in this study, which comprises of two experimental frameworks: (1)Invariance via Causal Data Augmentation and (2) Invariance via Causal Representation Learning.
@@ 3.1 Invariance via Causal Data Augmentation
Here, we investigate causal data augmentation as a strategy for improving sentiment classification in low-resource African languages through representation invariance. The approach is grounded in the principles of causal representation learning: if the sentiment label $y$ depends on the semantic content of a tweet $x$, then augmentations that perturb only non-causal linguistic features (e.g., surface form, style, or syntax) should preserve the conditional probability distribution $P(y|x)$.
@@
We used the Yoruba and Igbo subsets of the NaijaSenti corpus[7], a multilingual sentiment twitter dataset covering four Nigerian languages: Yoruba, Igbo, Hausa, and Nigerian Pidgin. Each instance is a tweet–label pair $(x_{i},y_{i})$, where $x_{i}$ represents the tweet and $y_{i}\in\{\text{positive},\text{negative},\text{neutral}\}$ is its sentiment label.
3.1.1 Experiment Procedure
Notation.
Let $\mathcal{X}$ denote the input text space and $\mathcal{Y}$ the sentiment label space. Let $\mathcal{D}_{1}=\{(x_{i},y_{i})\}_{i=1}^{N}$ be the original sentiment dataset sampled from a joint distribution $P(X,Y)$, where $x_{i}\in\mathcal{X}$ and $y_{i}\in\mathcal{Y}$. Let $f:\mathcal{X}\rightarrow\mathcal{X}^{\prime}$ be a paraphrasing function producing $x^{\prime}=f(x)$. We denote the paraphrased dataset as $\mathcal{D}_{2}=\{(x^{\prime}_{i},y_{i})\}_{i=1}^{N}$, and the combination of the paraphrased and original dataset as $\mathcal{D}_{\text{comb}}=\mathcal{D}_{1}\cup\mathcal{D}_{2}$.
Data Augmentation via Paraphrasing.
To simulate controlled domain shifts while preserving label semantics, we construct our paraphrased dataset $\mathcal{D}_{2}$ using a sentiment-preserving augmentation function $f:\mathcal{X}\rightarrow\mathcal{X}^{\prime}$ such that $P(Y\mid X^{\prime})=P(Y\mid X)$. The paraphrases are generated using the GPT-4o-mini model [18] with a carefully designed prompt that explicitly enforces sentiment and semantic preservation while introducing lexical and syntactic variation. The prompt template is shown below:
@@ 3.1.1 Experiment Procedure
“You have been provided with a {language} tweet. Paraphrase ‘{tweet}’ in a way that preserves its sentiment and meaning but changes its wording. Return only the paraphrased tweet.”
This process yields a paraphrased dataset $\mathcal{D}_{2}$ that is semantically aligned with the original dataset $\mathcal{D}_{1}$. We observed that GPT-4o-mini preserves sentiment polarity while introducing lexical and syntactic diversity which retains the underlying sentiment and lies in the same semantic neighborhood. The illustration in Table 2 are examples of original tweets and their corresponding paraphrases (‘Y” denotes Yoruba, ‘I” denotes Igbo).
Original Tweet ($X$) Paraphrased Tweet ($X^{\prime}$) (GPT-4o-mini) Label ($Y$)
Y: “Oúnjẹ yìí dára púpọ̀, mo fẹ́ràn rẹ̀ gan-an” “Oúnjẹ náà dùn púpọ̀, ó wù mí gidigidi” positive
I: “Ahịa a jọrọ njọ, ọ gaghị m azụta ya ọzọ” “Ngwa ahịa a adịghị mma, agaghị m alọghachikwuru ya” negative
Y: “Ayé ilá ilá ikàn wọ ẹ̀wù ẹ̀jẹ̀ sọ ọrún ọ̀pá gbédìí fáyé” “Ayé ilá ilá ikàn wọ aṣọ ẹ̀jẹ̀ sọ ọrún ọ̀pá gbédìí fáyé” negative
Table 2: Examples of sentiment-preserving paraphrases generated by GPT-4o-mini for Yoruba and Igbo tweets.
One of the main challenges in guiding LLMs to generate diverse variants while maintaining key information arises for proverbial and sarcastic tweets, where sentiment is often encoded through more nuanced expressions. In such cases, the model’s ability to retain variations of the original sentiment without diluting the underlying proverbial meaning or sarcasm is limited. To mitigate this issue, we employ a human annotation and correction mechanism for a subset of tweets exhibiting these challenges.
Our resulting output exhibits a shift in the marginal input distribution:
$\begin{split}P_{\mathcal{D}_{2}}(X^{\prime})&\neq P_{\mathcal{D}_{1}}(X),\\ P_{\mathcal{D}_{2}}(Y\mid X^{\prime})&\approx P_{\mathcal{D}_{1}}(Y\mid X).\end{split}$ (1)
Figure 2: Flow of $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, and $\mathcal{D}_{\text{comb}}$
For each language subset, we train three logistic regression classifiers: Original-only, trained on $\mathcal{D}_{1}$; Paraphrased-only, trained on $\mathcal{D}_{2}$; and Combined, trained on $\mathcal{D}_{\text{comb}}$ with a balanced 50–50 mix (this 50:50 mixing strategy is used to avoid bias stemming from over-representation of either original or paraphrased samples during training). We choose logistic regression for its interpretability and well-understood optimization properties, allowing us to isolate the effects of data augmentation and distributional shift without confounding factors introduced by model complexity. To assess robustness under the introduced distribution shift, we adopt a cross-domain evaluation protocol. Each model is evaluated on $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, and $\mathcal{D}_{\text{comb}}$, enabling systematic comparison of in-domain and cross-domain generalization performance. See Figure 3.
@@
Each classifier $h_{\theta}$ is trained by minimizing the empirical cross-entropy loss:
$\hat{\mathcal{L}}(\theta)=\frac{1}{|\mathcal{D}_{\text{train}}|}\sum_{(x,y)\in\mathcal{D}_{\text{train}}}\ell\big(h_{\theta}(g_{\psi}(\phi(x))),y\big),$ (2)
where $\phi$ denotes text preprocessing (tokenization, normalization, stopword removal), and $g_{\psi}$ is a fixed sentence embedding function implemented using the paraphrase-multilingual-MiniLM-L12-v2 model [19]. Our model performance was measured using “accuracy” for classification tasks. This experimental design enables a principled empirical test of our causal invariance hypothesis, by evaluating whether models trained with label-preserving augmentations improves generalization under distributional shift in low-resource settings.
Figure 3: Flow of the experiment procedure used in Invariance via Causal Data Augmentation.